OpenStudy (ttp):

Which is the same as 4 cm? zero point four millimeters forty millimeters zero point zero zero four meters four hundred meters

OpenStudy (doc.brown):

\[0.4mm\times\frac{1cm}{10mm}=\]\[40mm\times\frac{1cm}{10mm}=\]\[0.004m\times\frac{100cm}{m}=\]\[400m\times\frac{100cm}{m}=\]

OpenStudy (anonymous):

K H D M D C M ( King Henry Died Monday Drinking Chocolate Milk) 4 Kilo Hexa deka m/l/g deci centi milli When you move left move the decimal place over one. When you move right, move the decimal to the right. 2 is correct. :)

OpenStudy (ttp):

Thanks! :)

OpenStudy (anonymous):

1 cm = 10 mm 4 cm = 40 mm